OpenStudy (anonymous):

if f(x)=3-x^2, find f(2+delta x) .... i dont know how to do these problems using delta so please explain

OpenStudy (anonymous):

wherever x is there in the function f(x) to find f(2+delta x) replace x with 2 + delta x) so the function for f(x) which was 3- x^2 becomes 3-(2+delta x)^2 now we expand this, 3-(2+delta x)(2+delta x) =3-(2*2+2*(delta x)+ (delta x )* 2+(delta x)*(delta x)) =3-(4+4*(delta x)+(delta x)^2) =-1-4(delta x)-(delta x)^2 If you assume (delta x to be small,i.e. (delta x) <<1 then we can neglect the (delta x)^2 term)

OpenStudy (anonymous):

why do you have to assume that delta x is small?

OpenStudy (anonymous):

Obviously u could get -1 from simply substituting 2 into f(x). This is some sort of exercise (as a part of others, I would think) to show you how u may get to a result by showing how a small change in x leads to the same result. Delta x need not be assumed to be small but delta x is traditionally taken to mean a small change in x.

OpenStudy (anonymous):

oh ok thank you

OpenStudy (anonymous):

U r welcome.